OpenStudy (anonymous):

find f''(pi/4) using the equation f(x)=csc(x). I got 0 but it is wrong. What am I doing wrong?

OpenStudy (zarkon):

(csc(x))'=-csc(x)cot(x)

OpenStudy (zarkon):

then take derivative again using product rule

OpenStudy (zarkon):

then plug in pi/4

OpenStudy (anonymous):

yea i did that and got 0 but its wrong :/

OpenStudy (zarkon):

why don't you show your work

OpenStudy (anonymous):

\[\text{Cot}\left[\frac{\pi }{4}\right]^2 \text{Csc}\left[\frac{\pi }{4}\right]+\text{Csc}\left[\frac{\pi }{4}\right]^3=3 \sqrt{2} \]

OpenStudy (anonymous):

second derivative: csc(x)(cot^2x+csc^2x)